OpenStudy (anonymous):

Help me ? The equation of a circle is shown below. (x -5)2 + (y + 2)2 = 64 The radius of the circle is _______ units.

ganeshie8 (ganeshie8):

compare it wid standard circle equation : \((x-h)^2 + (y-k)^2 = \color{red}{r}^2\)

ganeshie8 (ganeshie8):

\(\color{red}{r}\) is the radius

OpenStudy (anonymous):

i dont get how i solve this.

ganeshie8 (ganeshie8):

given equation :- \((x -5)^2 + (y + 2)^2 = 64 \) \((x -5)^2 + (y + 2)^2 = \color{red}{8}^2\)

ganeshie8 (ganeshie8):

comparing wid the standard equation, r = 8, so radius = 8

OpenStudy (anonymous):

Okay i kind of get it but its still confusing .

ganeshie8 (ganeshie8):

its okay, do 2-3 more problems like these and you will feel comfortable im sure :)

OpenStudy (anonymous):

Thank you :)

ganeshie8 (ganeshie8):

np :) you're wlcme !